Calculate the pH of 0.65 M NaCN. What is the conce
Answer #1

Ka = 4.9 x 10^-10

NaCN ----> Na+ + CN-

0.65          0         0 (Initial)

0.65 - y       y          y ( Final)

CN- + H2O <-----> HCN + OH-

Kb = Kw/Ka

Kb = 1 x 10^-14 / 4.9 x 10^-10 = 2 x 10^-5

Kb = [Na+] [CN-] / [NaCN]

2 x 10^-5 = y^2 / ( 0.65 - y)

y^2 - 1.3 x 10^-5 + 2 x 10^-5 y = 0

y = 0.0036 M

[OH-]=0.0036 M

pOH = - log(0.0036) = 2.44

pH = 14 - 2.44 =11.56

[HCN] = 0.0036 M